OpenStudy (asadkarim7):

If the first three terms of the expansion of (1+ax)^n in ascending powers of x are 1-4x+7x^2, Find n and a. If the first three terms of the expansion of (1+ax)^n in ascending powers of x are 1-4x+7x^2, Find n and a. @Mathematics

OpenStudy (asadkarim7):

Please Helppp !!

OpenStudy (anonymous):

By the binomial expansion, (1+ax)^n = 1+(nC1)ax+(nC2)(ax)^2...

OpenStudy (anonymous):

nC1 = n, nC2 = n(n-1)/2

OpenStudy (anonymous):

Plugging in the values from the question, na = -4 and a^2(n^2-n)/2 = 7

OpenStudy (asadkarim7):

HEy Moorsy Got it !! Thanks 4 ur help :)

OpenStudy (asadkarim7):

1 more question why u did n(n-1) /2 . why did u divide the whole tern with 2 ??

OpenStudy (asadkarim7):

term *